OpenStudy (anonymous):

You have three $1 bills, four $5 bills, and two $10 bills in your wallet. You select a bill at random. Without replacing the bill, you choose a second bill at random. Find P($10 then $1).

OpenStudy (anonymous):

1/12

OpenStudy (hoblos):

P($10 then $1) = 2/9 * 3/8 = 6/72 = 1/12